### v2

# Jiafei Duan 1    Wentao Yuan 1∗    Wilbert Pumacay 2    Yi Ru Wang 1 Kiana Ehsani 3    Dieter Fox 1,4   Ranjay Krishna 1,3

Abstract
Large-scale endeavors like RT-1[1] and widespread community efforts such as Open-X-Embodiment [2] have contributed to growing the scale of robot demonstration data. However, there is still an opportunity to improve the quality, quantity, and diversity of robot demonstration data. Although vision-language models have been shown to automatically generate demonstration data, their utility has been limited to environments with privileged state information, they require hand-designed skills, and are limited to interactions with few object instances. We propose Manipulate-Anything, a scalable automated generation method for real-world robotic manipulation. Unlike prior work, our method can operate in real-world environments without any privileged state information, hand-designed skills, and can manipulate any static object. We evaluate our method using two setups. First, Manipulate-Anything successfully generates trajectories for all $5$ real-world and $12$ simulation tasks, significantly outperforming existing methods like VoxPoser. Second, Manipulate-Anything’s demonstrations can train more robust behavior cloning policies than training with human demonstrations, or from data generated by VoxPoser [3] and Code-As-Policies [4]. We believe Manipulate-Anything can be the scalable method for both generating data for robotics and solving novel tasks in a zero-shot setting.
Keywords: Zero-shot manipulation, multimodal language models, multiview state verification, robot skill generation, behavior cloning, robotic manipulation
1 Introduction
The success of modern machine learning systems fundamentally relies on the quantity [5, 6, 7, 8, 9, 10], quality [11, 12, 13, 14, 15], and diversity [16, 17, 18, 19, 20] of the data they are trained on. The availability of large-scale internet data made possible significant advances in vision and language [21, 22, 23]. However, the dearth of data has prevented similar advancements in robotics. Human demonstration collection methods do not scale to sufficient quantity or diversity. Projects like RT-1 [1] demonstrated the utility of high-quality human data collected over 17 months. Others have developed low-cost hardware for data collection [24, 25, 26]. However, all these procedures require expensive human collection. In an effort to diversify demonstration data, Open X-Embodiment project collected 1 million trajectories collected through a participatory effort by 34 research labs [2]. Despite the wide-spread effort, the dataset only contains 20 tasks.
Automated data collection methods do not scale to sufficient diversity. With the advent of vision-language models (VLMs), the robotics community has been abuz with new systems that leverage VLMs to guide robotic behavior [27, 4, 28, 29, 3, 30, 31]. In these systems, VLMs decompose tasks into language plans [27, 4] or generate code to execute predefined skills [32, 3]. Though successful in simulation, these methods underperform in the real world [32, 3]. Some methods rely on privileged state information only available in simulation [28, 30], require hand-designed skills [29]. or are also limited to manipulating a fixed set of object instances with known geometric shape [3, 32].
Figure 1: Manipulate-Anything is an automated method for robot manipulation in real world environments. Unlike prior methods, it doesn’t require priviledged state information, hand-designed skills, or limited to manipulating a fixed number of object instances. It can guide a robot to accomplish a diverse set of unseen tasks, manipulating diverse objects. Furthermore, the generated data enables training behavior cloning policies that outperform training with human demonstrations.
We propose Manipulate-Anything  a scalable automated demonstration generation method for real-world robotic manipulation. Manipulate-Anything produces high quality data, at large-quantities (if needed), and can manipulate a diverse set of objects to perform a diverse set of tasks. When placed in a real world environment and given a task (e.g., “open the top drawer” in Figure 2), Manipulate-Anything effectively leverages VLMs to guide a robotic arm to complete the task. Unlike prior methods, it doesn’t need privileged state information, hand-designed skills, or limited to specific object instances. Not relying on privileged information makes Manipulate-Anything environment-agnostic. Thus it can easily generalized to the real world. Manipulate-Anything plans a sequence of sub-goals and generates actions to excute the sub-goals. It can verify whether the robot succeeded in the sub-goal using a verifier and re-plan from the current state if needed. This error recovery enables mistake identification, re-planning, and recovering from failure. It also injects recovery behavior into the collected demonstrations. We further enhanced the VLMs’ capabilities by incorporating reasoning from multi-viewpoints, significantly improving performance.
We showcase the utility of Manipulate-Anything through two evaluation setups. First, we show that it can be prompted with a novel, never-before-seen task and complete it in a zero-shot manner. We quantitatively evaluate across $5$ real-world and $12$ RLBench [33] simulation tasks and demonstrate capabilities across many real-world everyday tasks (refer to supplementary). Our method significantly outperforms VoxPoser [3] in 9/12 simulation tasks for zero-shot evaluation. It also generalizes to tasks where VoxPoser completely fails because of its limitation to specific object instances. Furthermore, we demonstrated that our approach can solve real-world manipulation tasks in a zero-shot manner, achieving a task-averaged success rate of 36%. Second, we show that Manipulate-Anything can generate useful training data for a behavior cloning policy. We compare Manipulate-Anything generated data against ground truth human demonstrations as well as against data from VoxPoser[3] and Code-As-Policies [4]. Surprisingly, policies trained on our data outperforms even human data on 5 out of 12 tasks and performs on par for 3 more. Meanwhile, the baselines are unable to generate the training data for some of tasks. Manipulate-Anything demonstrates the broad possibility of large-scale deployment of robots across unstructured real-world environments. It also highlights its utility as a training data generator, aiding in the crucial goal of scaling up robot demonstration data.
2 Related work
Manipulate-Anything enables scaling of robotic manipulation data using VLMs. As such, we review recent efforts in 1) scaling manipulation data, and 2) applications of VLMs in robotics.
Scaling manipulation data. When deploying vision and language-based control policies for real-world applications, a significant challenge revolves around acquiring data. Traditionally, a convenient avenue to collect such trajectories is through human annotations for action (i.e. through teleoperation) and language labeling [34, 35, 36], however, this approach is limited to scale. To address this limitation and achieve autonomous scalability, prior works employ vision-language models or procedurally generate language annotations in simulated environments [30, 37, 38]. For action labels, strategies range from random exploration to learned policies [39]. While human egocentric videos are relevant, they lack action labels and require cross-embodiment transfer [40]. Another strategy involves model-based policies, such as task and motion planning (TAMP) [41]. Our approach extends these methods by incorporating common-sense knowledge from large language models (LLMs) and vision language models (VLMs), by providing a framework which combines the strengths of VLMs, object pose prediction, and dynamic retry to synthesize demonstrations in simulated and real environments.
Language models for robotics. In the field of robotics, large language models have found diverse applications, including policy learning [42], task and motion planning [43, 44], log summarization [45], policy program synthesis [4], and optimization program generation [27]. Previous research has also explored the physical grounding capabilities of these models [3, 32], while ongoing work investigates their integration with task and motion planners to create expert demonstrations [30]. [36] attempted to collect extensive real-world interaction data, with short-horizon trajectories. [46, 47] proposed a key-point based visual prompting method for real-world manipulation, through predicting affordances and corresponding motions. Our work complements the existing line of works, by leveraging the high-level planning capabilities of language models, scene understanding capabilities of vision language models, and action sampling, to enable synthesis of robot trajectories, which include language, vision, and robot state, given arbitrary tasks and environments.
3 Manipulate-Anything
Figure 2: Manipulate Anything Framework. The process begins by inputting a scene representation and a natural language task instruction into a VLM, which identifies objects and determines sub-goals. For each sub-goal, we provide multi-view images, verification conditions, and task goals to the action generation module, producing a task-specific grasp pose or action code. This leads to a temporary goal state, assessed by the sub-goal verification module for error recovery. Once all sub-goals are achieved, we filter the trajectories to obtain successful demonstrations for downstream policy training.
We propose Manipulate-Anything, a framework that solves everyday manipulation tasks conditioned on language. Under the hood, Manipulate-Anything leverages VLMs to decompose tasks into sub-goals, generates code for new skills or task-specific grasp pose, and verifies the success of each sub-goal (Figure 4). Note that due to the modularity aspect of our framework, Manipulate-Anything will continue to improve as the underlying VLMs continue to improve.
3.1 Task plan generation
Manipulate-Anything takes as input any task described by a free-form language instruction, $\mathbf{T}$ (e.g., ‘open the top drawer’). Creating robot trajectories that adheres to $\mathbf{T}$ is challenging due to its potential complexity and ambiguity, requiring a nuanced understanding of the current environment state. Given $\mathbf{T}$, and an image of the scene, we apply a VLM to first identify task-relevant objects in the scene, appending them to a list. Subsequently, We use a VLM to decompose the main task $\mathbf{T}$ into a series of discrete, smaller sub-goals, represented as $\mathbf{T}_{i}$, along with the corresponding verification conditions $v_{i}$, where $i$ ranges from 1 to $n$. For the above task, sub-goals include ‘grasp the drawer handle’ or ‘pull open the drawer’, and verification conditions are ‘did the robot grasp the handle?’ or ‘is the drawer opened?’. This transforms the instruction $\mathbf{T}$ into a sequence of specific sub-goals $\{(\mathbf{T}_{1},v_{1}),(\mathbf{T}_{2},v_{2}),\dots,(\mathbf{T}_{n},v_{n})\}$. For each sub-goal, Manipulate-Anything generates desired actions (§ 3.2) and uses the corresponding verification condition for each sub-goal to validate whether the generated actions result in the successful completion of the sub-goal (§ 3.3). This verification step allows Manipulate-Anything to recover from mistakes and attempt again in the case of failure.
3.2 Action generation module
Given a sub-goal, the desired output from the action generation module is a sequence of low-level actions represented as a 6 DoF end-effector pose. The actions can be categorized into two sets: agent-centric or object-centric. Agent-centric actions modify the agent’s state; e.g., it can move the robot’s end-effector from the current state (e.g., “rotate $90^{\circ}$”). We feed the VLM with the current observation along with in-context learning technique to write a code to synthesize the desired motion. Unlike prior methods that use only language models to generate code [4], our approach utilizes VLMs to understand and reason about object locations and the scene, which helps to ground the generation in the current state of the scene. This advantage is demonstrated in the ablation studies in the Appendix.
Object-centric actions require manipulating a certain object (e.g., “grasp a knife”). We use an object-agnostic grasp prediction model [48]. The grasp model generates all the possible 6-DOF grasping poses in the scene. These poses are not conditioned on the objects and could contain errors. From the RGB-D image of the current state, we extract a raw 3D point cloud. The point cloud is sent to the grasp model, which predicts 6-DoF grasps placements across the scene. We then further filter the proposed candidate grasp pose using VLM with in-context learning and condition it on the given task (e.g., if the task is “grasp a knife”, the VLM will detect the handle of the knife). Lastly, we use the detected bounding box to filter and sample an ideal grasp pose.
A single view point might be insufficient to provide the VLM with enough information to perform the task (e.g., some views might be occluded by the robot arm). Therefore, for both agent-centric and objec-centric action generation, we render multiple viewpoints of the scene and query VLMs to choose an ideal viewpoint given the sub-task. For example, if the task is to open a drawer, the view in which the handle of the drawer is visible would be preferred. After the best view point is chosen, the grasping poses can be filtered limited to the poses visible in that view point or the code generation will be conditioned on that image. After the action is generated, a simple motion planner can be used to move the robot to the desired pose as shown in detailed in Fig. 3.
Figure 3: Action Generation Module. Manipulate-Anything enables generation of two types of actions: object-centric and agent centric. For object-centric actions which require manipulation of an object, we leverage a foundation grasp model to generate all suitable grasps. Next, we leverage a VLM to detect the object from mult-view frames, and along with the candidate grasp poses and target subgoal, query the VLM to select the best view point. We filter and select the optimal grasp for the sub-goal. For more agent-centric actions, the view-point selection process is the same, and the goal is to output code representing the change in pose of the end-effector from the current frame.
3.3 Sub-goal verification
To ensure that each sub-goal $T_{i}$ is executed correctly, we introduce a VLM-based verifier. After every action for each sub-goal are executed, we use the VLM to check if the end state matches the verifier condition $v_{i}$. Similar to the action generation module, we use multi-view VLM reasoning to find the optimal view, avoiding errors due to occlusion or ambiguity from a single viewpoint. If the verifier identifies failure, we re-attempt the action generation step from the current state. Otherwise, the next sub-goal $T_{i+1}$ is attempted. More details of the implementation is in the Appendix.
Figure 4: Manipulate-Anything is an open-vocabulary autonomous robot demonstration generation system. We show zero-shot demonstrations for 12 tasks in simulation 5 tasks in the real world.
4 Experiments
Our experiments are designed to address two questions: 1) Can Manipulate-Anything accurately solve a diverse set of tasks in a zero-shot manner? 2) Can data generated from Manipulate-Anything be used to train a robust policy?
Implementation details. We use both GPT-4V and Qwen-VL [49] as our VLM. We use GPT-4V for task decomposition, action generation, and verification. We use Qwen-VL to detect and extract object information. To ensure zero-shot execution within a reasonable budget, we limit the number of action steps in each trajectory to $50$ and the verification module allows a maximum of $30$ tries to accomplish a sub-goal. For the task plan generation, we follow the prompting structure adapted from ProgPrompt [27]. All prompts input into the VLM are accompanied by few-shot demonstrations [50]. Additionally, we provide three manually curated primitive action code snippets as examples to prompt the VLM for new action code generation. Full prompts are included in the Appendix. We use four viewpoints $\mathbf{M}_{4}=[front,wrist,left\_shoulder,right\_shoulder]$ for the simulation experiments, and re-render three viewpoints for the real-world experiments [51]. For better reasoning by the VLM, we use a resolution of $256\times 256$.
4.1 Zero-shot Performance in Simulation
We empirically study the zero-shot capability of Manipulate-Anything in solving 12 diverse tasks in simulation. Our simulation experiments are reported to ensure reproducibility and provide a benchmark for future methods.
Table 1: Task-averaged success rate % for zero-shot evaluation. Manipulate-Anythingoutperformed other baselines in 9 out of 12 simulation tasks from RLBench [33]. Each task was evaluated over 3 seeds to obtain the task-averaged success rate and standard deviations.
Method Put_block Play_jenga Open_jar Close_box Open_box Pickup_cup
VoxPoser [3] 70.7±$\pm 2.31$ $\pm 0.00$ $\pm 0.00$ $\pm 0.00$ $\pm 0.00$ 26.7±$\pm 14.00$
CAP [4] 84.00±$\pm 16.00$ $\pm 0.00$ $\pm 0.00$ $\pm 0.00$ $\pm 0.00$ 14.67±$\pm 4.62$
MA (Ours) 96.00±$\pm 4.00$ 77.33±$\pm 6.11$ 80.00±$\pm 4.00$ 33.33±$\pm 12.86$ 29.00±$\pm 10.07$ 82.67±$\pm 14.04$
Method Take_umbrella Sort_mustard Open_wine Lamp_on Put_knife Pick_&_lift
VoxPoser[3] 33.33±$\pm 8.33$ 96.0±$\pm 6.93$ 8.00±$\pm 4.00$ 57.3±$\pm 12.22$ 92.00±$\pm 4.00$ 96.00±$\pm 0.00$
CAP[4] $\pm 4.00$ $\pm 0.00$ $\pm 0.00$ 64.00±$\pm 6.93$ 14.67±$\pm 8.33$ 100.00±$\pm 0.00$
MA (Ours) 61.33±$\pm 20.13$ 64.00±$\pm 6.93$ 42.00±$\pm 4.00$ 69.33±$\pm 6.11$ 52.00±$\pm 10.58$ 84.00±$\pm 6.93$
Environment and tasks. The simulation is set up in CoppeliaSim and interfaced through PyRep. All simulation experiments use a Franka Panda robot with a parallel gripper. Input observations are captured from four RGB-D cameras positioned around a tabletop setting. We use RLBench [33], a robot learning benchmark with diverse tasks conditioned on language and provided success conditions. We sample $12$ tasks from RLBench, covering a diverse range of action primitives, task horizons, and object position perturbations. Each action can be represented as a way-point, and the trajectories are computed and executed via a motion planner using the Open Motion Planning Library[52].
Baselines. We compare against two state-of-the-art zero-shot data generation approaches: Code-as-Policies (CAP) [4] and VoxPoser [3]. CAP uses language models to generate executable programs that call hand-crafted primitive actions. VoxPoser [3] builds a 3D voxel map of value functions for predicting waypoints. We provide both CAP and VoxPoser with ground truth simulation state information of the target object’s asset names or positions.
Results: Manipulate-Anything can generate successful trajectories for all $12$ tasks while VoxPoser and CAP cover only $8$ and $6$ tasks, respectively (Table 1). Without the privileged state information, the baselines would not succeed on any of the $12$ tasks. Manipulate-Anything outperforms the baselines in $9$ out of the $12$ tasks. The three tasks where our method achieves lower performance require fine-grained manipulation of objects, which are the hardest task without the privileged state information used by baselines. VoxPoser fails in the tasks that require moving the arm beyond 4-DoF. Manipulate-Anything outperforms the strongest baseline, VoxPoser, by an average task-averaged margin of up to $25\%$.
4.2 Behavior cloning with demonstrations from Manipulate-Anything
Next, we analyze the quality of the generated data by comparing the success rates of behavior cloning models trained with the data. Zero-shot methods like Manipulate-Anything are computationally expensive but hold the potential to generate useful training data. To evaluate the quality and effectiveness of the generated training data, we use the methods described in the previous section to generate data for each task. We also compare performance against a model trained on human-generated demonstrations across the $12$ tasks. We use the data to train behavior cloning policies.
Table 2: Behavior Cloning with different generated data. The behavior cloning policy trained on the data generated by Manipulate-Anything provides the best performance on 10 out of 12 tasks compared to the other autonomous data generation baselines. We report the Success Rate % for behaviour cloning policies trained with data generated from VoxPoser [3] and Code as Policies [4] in comparison. Note that the RLBench[33] baseline uses human expert demonstrations and is considered an upper bound for behavior cloning.
Generated data Put_block Play_jenga Open_jar Close_box Open_box Pickup_cup
VoxPoser[3] 2.67±$\pm 2.31$ $\pm 4.00$
CAP[4] 6.67±$\pm 2.31$ 14.67±$\pm 12.86$
MA (Ours) 85.33±$\pm 10.07$ 81.33±$\pm 2.31$ 21.33±$\pm 10.07$ 42.67±$\pm 8.33$ 30.67±$\pm 11.55$ 54.00±$\pm 12.49$
RLBench[33] 20.00±$\pm 18.33$ 81.33±$\pm 9.24$ 58.67±$\pm 45.49$ 68.00±$\pm 24.98$ 14.67±$\pm 6.11$ 54.67±$\pm 23.09$
Generated data Take_umbrella Sort_mustard Open_wine Lamp_on Put_knife Pick_&_lift
VoxPoser[3] $\pm 4.00$ $\pm 0.00$ 1.33±$\pm 2.31$ 5.33±$\pm 4.62$ 1.33±$\pm 2.31$ 5.67±$\pm 1.64$
CAP[4] 13.33±$\pm 10.06$ 8.00±$\pm 16.00$ 9.33±$\pm 6.11$ 46.67±$\pm 2.31$
MA (Ours) 84.00±$\pm 6.93$ 53.33±$\pm 6.11$ 86.67±$\pm 6.11$ 89.33±$\pm 6.11$ 8.00±$\pm 4.00$ 33.33±$\pm 2.31$
RLBench[33] 58.67±$\pm 50.80$ 53.33±$\pm 34.02$ 86.67±$\pm 12.86$ 84.00±$\pm 13.86$ 30.67±$\pm 10.07$ 62.67±$\pm 9.24$
Data generation details. We generate 10 successful demonstrations per task. We use the system’s success condition to filter for successful demonstrations. Each of the demonstrations consist of a language instruction, RGB-D frames for the trajectory, and waypoints represented as 6 DoF gripper poses and states. For the tasks that the baselines were unable to generate any successful demonstrations, we patched the missing training data with RLBench system-generated demonstrations.
Training and evaluation protocol. Using the generated demonstrations, we train a Perceiver-Actor (PerAct) model, which is a transformer-based robotic manipulation behavior cloning model [34]. The model expects tokenized voxel grids and language instructions as inputs and predicts discretized voxel grid 6 DoF poses and gripper states. For all the generated training datasets, we train a multi-task PerAct policy with a batch size of $4$ for $30$k iterations. To ensure consistent evaluation, we generate one set of testing environments with RLBench. We evaluate the last checkpoint from each of the trained policies. Each policy is evaluated for $25$ episodes across each task using $3$ different seeds. We measure the success rate based on the simulation-defined success condition.
Results: Policies trained using Manipulate-Anything data perform similarily to policies trained using human demonstrations ($p=0.973$) (Table 2). Training on either Manipulate-Anything or on human demonstrations results in a performance difference of a mere $0.27$% across all tasks. Furthermore, models trained on data from the baselines exhibit a statistically lower performance ($p\leq 0.01$ for both VoxPoser and CAP). One of the main factors potentially contributing to the differences in the performance could be that Manipulate-Anything generates diverse expert trajectories that are preferable to humans. This can be seen in Fig. 5, which shows the action distribution of the generated data by different methods for the same given tasks. Additionally, our generated data recorded the lowest Chamfer Distance (CD) of 0.056 with human-generated demonstrations data. We also observed that the policy trained on MA data achieves a lower standard deviation of $3.39$ across all tasks compared to zero-shot performance of $8.48$. This suggests the benefits of training over generated data instead of relying solely on zero-shot deployment.
Figure 5: Action Distribution for Generated Data: We compare the action distribution of data generated by various methods against human-generated demonstrations via RLBench on the same set of tasks. We observed a high similarity between the distribution of our generated data and thehuman-generated data. This is further supported by the computed CD between our methods and the RLBench data, which yields the lowest (CD=0.056).
4.3 Real-world experiments
Finally, we evaluate Manipulate-Anything in the real world. We also automatically generate real-world demonstrations for training PerAct.
Environment and tasks. We employ a Franka Panda manipulator equipped with a parallel gripper. We use a front-facing Kinect 2 RGB-D camera. To generate multi-view inputs for the Manipulate-Anything framework, we re-render virtual viewpoints from the generated point cloud, similar to prior work [51]. We select 5 representative real world tasks: open_jar, sort_objects, correct_dices, open_drawer, and on_lamp, all conditioned on language instructions. We evaluate each task for $10$ episodes, with varying object poses across $3$ trials of evaluation.
Data generation details. We used Manipulate-Anything to generate $6$ demonstrations for each task and manually perform scene resets when failures occur. We train a similar multi-task PerAct for $120$k iterations and evaluate the trained policies in a manner similar to the zero-shot experiments.
Table 3: Real-world Results. The model trained on the data generated by our model in the real world (no expert in the loop) demonstrates on par results with the model trained on human expert collected data. We present a comparison of success rates for task completion in a zero-shot manner (Code as Policies [4] and Manipulate-Anything), and using trained policies from Manipulate-Anything data and human expert data.
Open_drawer Sort_object On_lamp Open_jar Correct_dice
CAP (0-shot) $00.00\pm 0$ 0.00 $13.33\pm 5.77$ $00.00\pm 0$ 0.00 $6.67\pm 5.77$ $6.67\pm 5.77$
MA (0-shot) 36.67±$\pm 5.77$ 60.00±$\pm 10.00$ 26.67±$\pm 11.55$ 40.00±$\pm 10.00$ 53.33±$\pm 5.77$
PerAct (MA data) $50.00\pm 0.00$ $33.33\pm 5.77$ $46.67\pm 5.77$ $56.67\pm 5.77$ $60.00\pm 0.00$
PerAct (Human data) 53.33±$\pm 11.55$ 36.67±$\pm 5.77$ 60.00±$\pm 0.00$ 76.67±$\pm 5.77$ 80.00±$\pm 10.00$
Results: Manipulate-Anything is able to generate successful demonstrations for each of the $5$ real world tasks. Even for the worst-performing task, Manipulate-Anything achieves a success rate of more than $25\%$. Our approach outperforms CAP by 38%. Consistent with the simulation results, training with the data generated by Manipulate-Anything produces a more robust policy compared to performing zero-shot. Additionally, in 4 out of 5 tasks, the trained policies perform better than the zero-shot approach. The policy underperforms on the sort_object task, because it requires longer-horizon memory —a known limitation pointed out in PerAct [34].
Figure 6: Scaling experiment. Scaling effect of model performance with increasing training demonstrations.
4.4 Ablations
For effective real-world deployment of Manipulate-Anything, it’s crucial that the collected data supports scaling of robotics transformers and offers diverse skills and interacted objects. We conducted an ablation study to evaluate the quality of Manipulate-Anything-generated data for scaling and its generalization to language instruction changes. For scaling, we generated behavior cloning data, ranging from 1 to 100 training demonstrations from RLBench and Manipulate-Anything for a single task, and trained a PerAct policy. For generalization, we varied the sort_mustard task with different language instructions and target objects. We compared our approach to VoxPoser to assess robustness to object and language instruction changes. Further implementation details are in the supplementary materials. Result: Our scaling experiments demonstrate that generating more training data via Manipulate-Anything improves PerAct policy performance (Fig. 6). The data from our approach shows a better rate of change with a slope of 0.503 for a linear fit, compared to 0.197 for RLBench-generated data. Additionally, Manipulate-Anything data is more generalizable and robust to language instruction changes, outperforming VoxPoser in task success across language and object variations. Detailed results in the appendix.
5 Discussion
Limitations. Manipulate-Anything relies on the availability of VLMs. While this can pose a dependence on the foundational models with the rise of the open VLMs, we believe this issue will be addressed soon. Future work. With the enhancements of large foundational models, Manipulate-Anything, due to its modularity, will continue to grow and scale up to more complex tasks. Conclusion. Manipulate-Anything is a scalable environment-agnostic approach for generating 0-shot demonstration for robotic tasks without the use of privileged environment information. Manipulate-Anything uses VLMs to do high level planning and scene understanding and is capable of error recovery. This enables high quality data generation for behavior cloning that can achieve better performance that using human data.
6 Acknowledgement
Jiafei Duan is supported by the Agency for Science, Technology and Research (A*STAR) National Science Fellowship. Wilbert Pumacay is supported by grant 234-2015-FONDECYT from Cienciactiva of the National Council for Science, Technology and Technological Innovation (CONCYTEC-PERU). This project is partially supported by Amazon Science. We would also like to thank Winson Han from AI2 for helping with the figure and icon design.